# Counterfactual Effect Generalization: A Combinatorial Definition

### v4

@@ 3 Sampling Treatment-Background Assignments
That’s because, for a fixed gap index $k$, it takes $2^{t}\times k(\bmod 2n-1)=k$ time for $u_{k}$ to get back to its original position in this case [Diaconis et al., 1983], and for all background factors to be enumerated222as result of the recursion $(2(2...(2k(\bmod 2n-1))...)))$ for each gap position $k$.. This can be restated as
$2^{t}(\bmod 2n-1)=2^{t}\ (\bmod\ q-1)=1,$ (10)
for a initial position $k=1$.
The impact of the observation of increasing backgrounds $Z/X$ on sample effects (and background enumeration) depends on the specific extension between posets $G=\{G(X),G(U)\}$, which we exemplified in Example.1, described as the partition of $U$ induced by sample gaps, Eq.(2), and now describe stochastically with a multinomial distribution, $p=\{p_{1},p_{2},...,p_{m}\}$. Each distributions (and the linear extensions they imply) is associated with a number of permutations of $U$ they can generate under a fixed $X$. This was formulated in Eq.(2), but can be made more precise by fixing an initial permutation $\pi_{0}$ and counting the number of inversions, or pairs of elements that are in the reverse order, from $\pi_{0}$. The expected number of permutation inversions, $I(\pi_{0},t)$, generated by a fixed distribution $p$ and initial background $\pi_{0}$ (Appendix.A) after time $t$ is
${\rm I\kern-1.99997ptE}[\,I(t\,|\,p,\pi_{0})\,]$ $=\frac{1}{2}\times\binom{m}{2}\times\Big{(}1-\Big{[}p_{1}^{2}+p_{2}^{2}+...+p_{m}^{2}\Big{]}^{t}\Big{)}$ (11)
where $\binom{m}{2}$ denotes the number of effect observations in a single square. For the case of $p=\{0.5,0.5\}$ the equation reduces to
${\rm I\kern-1.99997ptE}[\,I(t\,|\,\pi_{0})\,]$ $=\binom{m}{2}\times\Big{(}1-\frac{1}{2^{t}}\Big{)}$
@@
which underlies the known cut-off limit for markov chains  [Diaconis, 1996, Bayer and Diaconis, 1992]. As consequence, under random backgrounds, effect generalizability requires $2^{t}$ effect samples, according to Definition.5, in a fixed or fast $U-X$ system. Another trivial example has ordinal probability mass $p=\{1,0,0,0,...\}$ concentrated in a single gap, which, as expected, generates no inversions. This leads to the limit $|\Pi_{n}|=0$ which describe samples with no EV non-increasing sequences. An unobserved sample $U$ close to $p=\{1/q,1/q,1/q,1/q,...\}$ ($m\geq q$) is capable of asymptotically enumerating $|\Pi_{n}|=q!$ ordering of unobserved causes in least time $t$. The number of permutations, for a given $t$, is decreasing as we depart the second (random $U$) and approach the first (deterministic $p=\{1,0,0,0,...\}$). For intermediary and more typical scenarios, what matters is the particular distribution $p$ of confounders across a sample’s gaps.
Other than Eq.(11), there are many ways to show that any departure from uniformity will only decrease the number of EV-increasing sequences in the sample. Consider first the case of a stationary gap probability $p_{t}$. A new background is generated every $m/q$ time. Since $m$ is known and fixed, the $p_{t}$ where each and every factor effect is observed under the same background can be written as $p_{t}=(m/q,m/q,m/q,...)$ (the conditional uniform distribution) which can, in turn, be written as $p_{t}/m=(1/q,1/q,1/q,...)$. A simple approximation is $p/m=1/q+d_{j}$ where $\sum d_{j}=0$. Define then $s=\sum d^{2}$ as a ’phase’ between ${U,X}$ systems. This type of multinomial distribution approximation appear in the context of the birthday paradox  [Nunnikhoven, 1992, Gail et al., 1979] and reflects the recursive solution,
$S_{q,m}\Big{\{}\,p=[p_{1},...,p_{m-1}]\,\Big{\}}$ $=m!\sum_{X_{(1)}<X_{(2)}<...<X_{(m)}}p_{i}\times...\times p_{m}$ (12)
$=S_{q,m-1}\{[p_{1},...,p_{m-1}]\}+p_{m}S_{q-1,m-1}\{[p_{1},...,p_{m-1}]\}$ (13)
where the first equation summing over all subsets of $U$ such that $X_{(}1)<X_{(}2)<...<X_{(}m)$ and the second with initial conditions $S_{1,j}=\sum_{l=1}^{j}p_{i}$, and $S_{j,j}=p_{1}\times...\times p_{j}$ for $j=1,2,...,m$. Assuming a finite $q$, with each new sample ($m<q$), a further $p_{j}$ (rank) is observed, further approximating the (unknown) $p$ ($m\geq q$). The $s$ index is simply an indicator of changes in $U{-}X$ ’phase’ during this process, and will be useful when testing hypotheses about their relative changes.
For the case where all in-sample variables have the same ordinal frequencies - with their effect observations illustrated by the square diagram, Fig.1(b, lower-right) - $d$ describes departure from ideal $U{-}X$ effect background enumeration ($d$ is subquadratic and $q,m$ coprime), Definition.5. When $d\approx 0$, there is no ’phase’ between $X$ and $U$, and the sample treatment-background assignment mechanism is fixed. When $U$ and $p$ are fixed and $d>0$, effect backgrounds can be enumerated, due to its additive effect on $b(t)$, $\pi_{0}:b(t)=t+d\bmod q$ When $U$ is changing but $p$ is fixed, $d$, instead, has a multiplicative effect, $\pi_{0}:b(t)=d^{-t}\bmod q$, on enumeration and quickly becomes random333as result of the recursion $(d^{-1}(d^{-1}...(d^{-1}k(\bmod q-1))...)))$ for each gap position $k$. (an observation often employed by random number generators  [Knuth, 1997]). The relationship between $d$ and randomness is a deep, and somewhat unresolved, issue in mathematics  [Bayer and Diaconis, 1992, Diaconis and Fulman, 2023]. In the context of effect estimation and $U{-}X$ systems, this is a specially relevant case (i.e., of many time varying confounders), as it characterizes most real-world data and is where out-of-sample generalization is most useful. In this case, we want guarantees that any treatment-background assignment selected after $\pi_{0}$ could be chosen with the same probability (i.e., any of their orderings is as ’random’ as any other).
@@
Under changing backgrounds $Z/X$ and fixed $p$, effect treatment-background enumeration can become asymptotically ergodic, and thus EV-increasing with high-likelihood; for an initial effect background $\pi_{0}$, the sample sizes required for this condition depend on the treatment-assignment distribution $p$, and the number of inversions from $\pi_{0}$ the distribution $p$ can generate, Eq.(11) (Sect.4.2 Effect Background Randomization).
This leads to the proposed solutions here for causal effect estimation with increasing generalizability,
@@
Counterfactual background effect enumeration (Definition.6) is EV increasing in fast $U{-}X$ systems, and background randomization (Definition.5) is EV increasing in slow systems.
@@ 4 Related Work
Non-parametric tests (like permutation tests) can often be used only under assumptions of strict independence among factors and no unobserved factors. Consider, for example, a scenario where two sample units are observed (but not randomized), and we pair each unit receiving a first treatment $a$ with a different unit receiving $b$. To test the null hypothesis that the two treatments have no difference on any unit, we can permute treatment assignments within the pairs. The strategy relies on the assumption that the two units in the pair are surely exchangeable besides their treatment status. This assumption can be violated if the units are different in any unobserved way (even in high-dimensional samples it is hard for researchers to make the claim that there are no overlooked and relevant distinctions). Randomization plays a crucial role in alleviating the effects of unobserved confounding  [Zhang and Zhao, 2023, Kempthorne and Doerfler, 1969]. Randomized Experiments provide a solid foundation, and are still the golden rule in the Sciences, for making causal inferences about the effects of treatments on an outcome. Non-experimental samples can be seen as experimental samples where randomization did not occur perfectly. This leads to a scenario where treatment or control group are different, with units in one of them having factors that are predictive of (or are merely correlated with) the outcome, thus making treatment-control outcome differences invalid effect estimates. In these scenarios, standard non-parametric techniques becomes inapplicable. Despite these known shortcomings, non-parametric tests, specially permutation-based tests like the Shapley value  [Ribeiro, 2022a, Burkart and Huber, 2021], continue to be used in practice in non-i.i.d.and partially observed data, with little clarity over the resulting sample biases and generalizability.
The previous discussions illustrated the need to also permute (unobserved) effect backgrounds when estimating effects. According to our view, due to the inherent unobservability of those factors, the challenge for estimation then becomes how to establish in-sample conditions indicating when backgrounds (1) have not changed, or, (2) that they have changed enough such that any association between treatment and backgrounds of two effect observations or units nearly vanished. We suggested that a way to establish such conditions when $|\Pi_{n}^{*}(Z/\{a\})|\gg 0$ is to seek assurances of randomness from established theory, such as from Ergodic and Combinatorial Randomization theory, as further discussed below. In respect to the generalization of effects, another suggestion was that there are advantages in carrying randomization jointly across all $X$ factors.
@@
Concerns about robustness and predictive performance can, in part, be seen as the motivation for more recent views of causality based on model-performance invariance  [Peters et al., 2016, Magliacane et al., 2017, Buehlmann, 2020]. This is typically the invariance of a black-box model’s performance - and not of an effect, or, difference in outcomes, in a population, as taken here. The unit of analysis here, $\Delta y$, is much simpler (being an observation and not an estimate). Furthermore, on the present view, it is not the minimum variance of effects that matters, but the minimum variance of effects in face of the maximum variance of extraneous factors. The approach here addresses the latter conditions, and shed light on when successive observation of effects can lead to increases in out-of-sample performance. Finally, we believe the widespread success of the counterfactual definition is largely due to its simplicity, and simplicity has been a central guideline in the present work.
@@
The issue of the Extenal Validity (EV) of effects has been addressed in  [Pearl and Bareinboim, 2014, 2011, Correa and Bareinboim, 2020]. Authors assume a known ’selection’ variable set $S$ that ’may represent all factors by which populations may differ or that may threaten the transport of conclusions between populations’ when training parametric models on non-experimental data. They then, assuming causal graphs for all other variables are known or inferred for both population, propose graphical conditions for EV based on Pearl’s graphical calculus. Assuming all this prior knowledge is known, the proposed conditions may give assurances on the transportability of ’observational findings’. Often those differences are not known. We take a more practical starting point, where every variable is potentially a relevant difference, no independence relations are assumed, and all results follow from in-sample effect differences and probabilities. We assume that subsets of ’population’ factors are observed ($X$) but another ($U$) is completely unobserved, as typical in supervised training tasks. We provide a quantifiable measure of EV (not a set of ideal graphical and independence conditions) for samples. In our view, the EV of effects is an uncertainty measure, Eq.(4), and not an isolated issue. At the same time, it is difficult to think of issues of external validity separate of issues of selection bias (which is the central issue in scientific research and Experimental Design)  [Cox, 2009, Montgomery, 2001]. The solution is formulated at the level of individual effect observations, non-parametrically, and in the counterfactual framework. We consider whether connections to Combinatorial Randomization might be helpful. The alternative combinatorial model proposed allow us to discuss critical issues - difficult to address, even if only theoretically, in the previous - like omitted variables, in-sample and out accuracy trade-offs, selection bias, prediction-interpretability tradeoffs, empirical results, and sample size requirements.
4.1 Randomization and Causal effects

### v5

@@ 3.1 Background Enumeration
For a fixed background $Z/X$ (and a constant sampling rate), sequences $b(t+1)=c_{1}\times b(t)+c_{2}\bmod q$ for $\nicefrac{{c_{1}}}{{q}}$ integral and $\text{gcd}(c_{2},q)=1$ are EV increasing and CF non-increasing, where $\text{gcd}(c_{2},q)=1$ indicate that $c_{2}$ and $q$ are co-primes.
Consider, for example, $c_{1}/q=2$. For a fixed gap index $k$, it takes $2^{t}\times k(\bmod 2n-1)=k$ time for $u_{k}$ to get back to its original position  [Diaconis et al., 1983], and for all background factors to be enumerated111as result of the recursion $(2(2...(2k(\bmod 2n-1))...)))$ for each gap position $k$.. This can be restated as
@@
for a initial position $k=1$. This condition is known to only ’circulate’ the set of factors, $b(t)$, and lead to no randomization  [Diaconis et al., 1983, Diaconis and Fulman, 2023]. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}<u_{4}\;\}$, for example, where $q=4$ and $c_{1}=2$, linear extensions of type $\mathbf{a}u_{1}u_{2}\mathbf{b}u_{3}u_{4}$ can generate EV-increasing sequences with a constant sampling rate, unlike $\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b}u_{4}$.
3.2 Background Randomization
We next consider the case of in-sync and slow sampling, $q\gg m$. Both in-sync and slow sampling can be seen as allowing the multiplicative and integral constant $c_{1}$ to take random values. In the in-sync case, $c_{1}$ takes approximately common $\Delta_{(k)}q$ values across all $k$, $0<k\leq m$. It then describes background-treatment assignments where the $q$ confounders are equally distributed across sample gaps, $\Delta_{(k)}q\times m\approx q$. In Example (1), we exemplified how sample effect estimation depends on the extensions between posets $G=\{G(X),G(U)\}$. We then described them as a partition of $U$ induced by sample gaps in Eq.(2). More specifically, we can describe this relationship with a multinomial sample gap distribution, $p=\{p_{1},p_{2},...,p_{m}\}$.
Let $p=\{p_{1},p_{2},...,p_{m}\}$ then be a sample gap background assignment. Consider a set of subsequent unobserved $U=Z\backslash X$ backgrounds, Definition (1). Generalizability requires effects to be valid across any $\Pi_{n}(Z/X)$, irrespective of $p$. Imagine we keep $p$ constant from time $t=0$ to $\delta t$ while allowing $\Pi_{n}(Z/X)$ to change. Can we say anything about whether the backgrounds $\pi_{0}$ and $\pi_{\delta t}$ are random-like (i.e., the realization of $U$ in the first is as ’separated’ to the second as independent samples of a uniform distribution), $\pi_{0}\perp\pi_{t}$? We can describe this relationship more precisely by fixing an initial permutation $\pi_{0}$ and counting the number of inversions, or pairs of elements that are in the reverse order, from $\pi_{0}$ at the subsequent time $t$. Fix $p$ and two backgrounds $\pi_{0},\pi_{1}\in\Pi_{n}(Z/X)$ separated by only one inversion, for example. The likelihood of reaching $\pi_{1}$ from $\pi_{0}$ after one time step is high in this case. With increasing inversions in $\pi_{0}$, there should be some period $\delta t$ where that likelihood becomes comparable to that of a uniform distribution, where the expected number of inversions generated at each time is $\binom{m}{2}$. The expected number of permutation inversions, $I(\pi_{0},t)$, required for that condition (’randomization’) with a fixed distribution $p$, initial background $\pi_{0}$, and under i.i.d. sampling  [Fulman, 1998] is
${\rm I\kern-1.99997ptE}[\,I(t\,|\,p,\pi_{0})\,]$ $=\frac{1}{2}\times\binom{m}{2}\times\Big{(}1-\Big{[}p_{1}^{2}+p_{2}^{2}+...+p_{m}^{2}\Big{]}^{t}\Big{)},$ (11)
after time $t$. With a $p$ close to $\{1/q,1/q,1/q,...\}$, randomization happens quickly, as the term within the right-most parenthesis in Eq.(11) converges quickly to $1$, with an increasing $t$. The choice of a background $\pi_{1}$ from $\pi_{0}$ is then as likely as any other possible background. For the case of $p=\{0.5,0.5\}$ the equation reduces to
${\rm I\kern-1.99997ptE}[\,I(t\,|\,\pi_{0})\,]$ $=\binom{m}{2}\times\Big{(}1-\frac{1}{2^{t}}\Big{)},$ (12)
@@ 3.2 Background Randomization
which underlies the known cut-off limit for markov chains  [Diaconis, 1996, Bayer and Diaconis, 1992]. As consequence, under random backgrounds (and where $U{-}X$ sampling is i.i.d.), effect generalizability requires $2^{t}$ effect samples. This is a stochastic generalization of Definition (5) and the previous case of $c_{1}=2$ on fast $U-X$ systems, $\nicefrac{{\Delta_{(k)}}}{{q}}\sim Bin(q,0.5)^{-1}$. Another trivial example has probability mass $p=\{1,0,0,...\}$ concentrated in a single gap, which, as expected, generates no inversions. This leads to the limit $|\Pi_{n}|=0$ which describes samples with no EV increasing sequences. A sample with $p$ close to $p=\{1/q,1/q,1/q,...\}$ ($q\geq m$) is capable of asymptotically enumerating $|\Pi_{n}|=q!$ orders of unobserved causes in the least time $t$, Eq.(2). This condition is associated with $U{-}X$ independence, and, departures from this condition with more restrictive requirements on sampling times or sample sizes. The number of EV-increasing permutations, for a given $t$, is decreasing as we depart the last limit (random $U$) and approach the previous (deterministic $p=\{1,0,0,0,...\}$). These cases are then associated with sample sizes of, respectively, order $\mathcal{O}(log_{2}t)$ and infinite for generalizable effect observations. For intermediary and more typical scenarios, what matters is the particular distribution $p$ of confounders across a sample’s gaps, Eq.(11). Notice that $\binom{m}{2}$ in Eq.(11) is the number of effects observation in a single square.
The relationship between inversions and randomness is a deep, and somewhat unresolved, issue in mathematics  [Bayer and Diaconis, 1992, Diaconis and Fulman, 2023]. With an assumption of stationarity of $p$ (Sect.4 Square Enumeration and Gap Randomization Tests), Eq.(11) suggests under what conditions (and time after $\pi_{0}$) we can say that any treatment-background assignment selected would be chosen with equal probability.
@@
Under non-fixed backgrounds $Z/X$ and stationary $p$ (i.i.d.), sampling sequences $b(t+1)=c_{1}\times b(t)+c_{2}\bmod q$ and effect treatment-background enumeration can become asymptotically random, and thus EV-increasing with high-likelihood. For an initial effect background $\pi_{0}$, sample sizes required for this condition depend on the treatment-assignment distribution $p$, and the number of inversions from $\pi_{0}$ the distribution $p$ can generate, Eq.(11).
Fig.2(b) illustrates this limit for the simulations in Sect.7 Experiments, which include samples with unobserved, correlated and non-stationary factors. The variation separation distance (y-axis) is a measure of randomness formulated as a distance to an uniform distribution  [Bayer and Diaconis, 1992, Diaconis, 1996]. The panel shows the distance between a fixed background $\pi_{0}$ and the subsequent in samples where $p$ is stationary (according to a Kendall-tau test, permutations and inversions are exhaustively enumerated and counted using the tests and procedures described next, distances where normalized to their per-factor range for display). The upper panel shows the same for other samples where $p$ is not stationary (according to Kendall-tau). These combinatorial considerations suggest limits where some guarantees on effect background randomization can be made. In systems with $q\gg m$, they guarantee that effects are observed in new backgrounds after each formulated period and, thus, that effects are observed under EV-increasing sampling sequences.
The proposed condition is only valid for a stationary $p$ and initial $\pi_{0}$. We associate EV-CF sequences, instead, with $m$ sampling sequences with the same properties but observed under $m$ in-sample cyclic permutations (i.e., with period $m\times\delta t$). The consequence of the later is that effects for all factors are observed across the same backgrounds for all in-sample factors $X$. These i.i.d. and permutation inversion conditions are further discussed in Sect.4 Square Enumeration and Gap Randomization Tests. In the context of effect estimation, Eq.(11) relates slow sampling to in-sync sampling (as visualized by squares) and the conditions under which they can lead to samples with generalizable effects. They implement the stipulated gap distributions in Eq.(8) and lead to the proposed solutions here for causal effect estimation with increasing generalizability,
@@
Counterfactual background effect enumeration, Definition (5), is EV increasing in fast $U{-}X$ systems, and background randomization, Definition (6), is EV increasing in in-sync and slow systems.
4 Square Enumeration and Gap Randomization Tests
Both the randomization conditions in the last section and Principle (1) are related to permutation inversions. Inversions appear in common tests for i.i.d. sampling (’stationarity’) in ordinal data, such as in Kendall-tau and Man-Whitney tests. Any pair of observations $(x_{i},y_{i})$ and $(x_{j},y_{j})$, where $i<j$, are said to be ’concordant’ if the sort order of $(x_{i},x_{j})$ and $(y_{i},y_{j})$ agrees: that is, if either both $x_{i}>x_{j}$ and $y_{i}>y_{j}$ holds or both $x_{i}<x_{j}$ and $y_{i}<y_{j}$; and be ’discordant’ otherwise. The number of discordant pairs between $x$ and $y$ is equal to the number of inversions that permutes the $y$-sequence into the same order of $x$. The statistic can also be seen as a generalised measure of a classifier’s separation power for more than two classes  [Hand and Till, 2001].
We now formalize the previous conditions for effect background randomization in non-experimental samples. They are fulfilled by any subsample $C$, $C\subset X$, passing the following statistical hypothesis tests over inversions,
[…]
We can also generalize the previous conditions to samples that are not sequentially sampled. This is done by enumerating squares for a specific sample unit $x_{0}$ using inter-unit combinatorial properties. In fully observed samples, $q=0$, square enumeration ensures the applicability of Principle (1). In samples with unobserved factors, $q>0$, enumeration with the additional conditions over gap inversions in Eq.(14) offer analogous advantages. Much like we can observe single factor differences in samples, and their effects, we can observe factor permutations and their effects, by combining across-unit differences and their effects carefully. The panel in Fig.2(d) illustrates this process for two factors ($m=2$). It illustrates the differences (circles) and intersections (gray), from $x_{0}$, that characterize each cell in a square of size $2$. We say that by observing differences $a$, $b$ and $ab$ from $x_{0}$, we ’observed’ a permutation for these factors. That’s because these differences - $(x_{0}{-}x_{1})=\{a\}$, $(x_{0}{-}x_{2})=\{b\}$ and $(x_{0}{-}x_{3})=\{ab\}$ allow us to recover the effects of permuting factors $\{a,b\}$ (conditional on $x_{0}$), which are not directly observed in the sample. Notice that not all sample pairs with differences $\{a\}$ and $\{ab\}$ would belong to the second square column, as there are also conditions on their intersection, $(x_{0}{-}x_{1})\cap(x_{0}{-}x_{3})=\{b\}$. It is the fulfillment of both combinatorial conditions that allow us to take them as (conditional) effect observations. The square full combinatorial structure generalizes this to $m>2$ and uses the $2^{m}$ $m$-way differences potentially available in samples. Further combinatorial, computational and statistical aspects of this enumeration process are described in Appendix.A,B,I.
Fig.1(a) summarizes the proposed solution for effect estimation, where square enumeration ensures the inversions required by Principle (1) for in-sample factors, and tests over their gaps, Eq.(14), for out-of-sample factors. For sets of units (and thus mutual effect observations) where these combinatorial and gap conditions hold simultaneously, variance in effects can become asymptotically indicative of the status of in-sample factors as causes, Principle (1), their effects, Eq.(3), and their effects generalizability, Eq.(4). Variance becomes, in resulting samples, a measurement of ’noise’ for individual effect observations that also takes causal principles into consideration  [F. Ribeiro et al., 2022].
@@ 5 Related Work
Non-parametric tests (like permutation tests) can often only be used with strict assumptions on the completeness of samples. The previous discussions suggested the need to permute (unobserved) effect backgrounds when estimating effects. According to our view, due to the frequent partial observability of samples, the challenge for estimation then becomes how to establish in-sample conditions indicating when backgrounds (1) have not changed, or, (2) that they have changed enough such that any association between treatment and backgrounds of two effect observations or units nearly vanished. We suggested that a way to establish such conditions, when the number of unobserved factor is large, is to seek assurances of randomness from established theory, such as from Ergodic and Combinatorial Randomization theory  [Diaconis, 1996, Diaconis and Fulman, 2023].
@@
Concerns about robustness and predictive performance can, in part, be seen as the motivation for more recent views of causality based on model-performance invariance  [Peters et al., 2016, Magliacane et al., 2017, Buehlmann, 2020]. This is typically the invariance of a black-box model’s performance - and not of an effect, or, difference in outcomes, in a population, as taken here. The unit of analysis here, $\Delta y$, is much simpler (being an observation and not an estimate). Furthermore, on the present view, it is not the minimum variance of effects that matters, but the minimum variance of effects in face of the maximum variance of extraneous factors. The approach here addresses the latter conditions. Finally, we believe the widespread success of the counterfactual definition is largely due to its simplicity, and simplicity has been a central guideline in the present work.
@@
The issue of the Extenal Validity (EV) of effects has been addressed in  [Pearl and Bareinboim, 2014, 2011, Correa and Bareinboim, 2020]. Authors assume a known ’selection’ variable set $S$ that ’may represent all factors by which populations may differ or that may threaten the transport of conclusions between populations’ when training parametric models on non-experimental data. They then, assuming causal graphs for all other variables are known or inferred for both population, propose graphical conditions for EV based on Pearl’s graphical calculus. Assuming all this prior knowledge is known, the proposed conditions may give assurances on the transportability of ’observational findings’ (parameter estimation). Often the previous differences are not known. We take a more practical starting point, where every variable is potentially a relevant difference, no independence relations are assumed, and all results follow from in-sample effect differences and probabilities. We assume that subsets of ’population’ factors are observed ($X$) but another set of factors ($U$) is completely unobserved, as typical in supervised training tasks. We provide a quantifiable measure of EV (not a set of ideal graphical and independence conditions) for samples. In our view, the EV of effects is an uncertainty measure, Eq.(4), and not an isolated issue. The solution is formulated at the level of individual effect observations, non-parametrically, and in the counterfactual framework. At the same time, it is difficult to think of issues of external validity separate of issues of selection bias (which is the central issue in scientific research and Experimental Design)  [Cox, 2009, Montgomery, 2001] and randomization. We consider whether connections to Combinatorial Randomization might be helpful. The alternative combinatorial model proposed allow us to discuss critical issues - difficult to address, even if only theoretically, in the previous - like in-sample and out accuracy trade-offs, selection bias, prediction-interpretability tradeoffs, empirical results, and sample size requirements.
5.1 Randomization and Parametric Inference